I have been trying for two days on this math. Please help me if you can.
Graph the linear inequality 2x-y=5

Answer by mananth(16949)

2x-y=5
y=2x-5
when x=0 y=-5 (0,-5)
when y=0,x=5/2 (5/2,0)
Plot these points and draw the line graph.
you area there